Greenhab Report – April 23rd

Crew 211 GreenHab Report – 23-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 2

Shade cloth: 80%

Average temperatures:

Low: 13°C

High: 41°C

Hours of supplemental light: 0

Daily water usage for crops: 3 gallons

Water in Blue Tank: 110 gallons

Time(s) of watering for crops: 16:00

Changes to crops: All plants look healthy. Flowers are visible on parsley. Romaine lettuce still grow well.

Narrative: Still hot today. Tomorrow I will run the fan on 3 between 10:00 to 18:00 as you recommended me yesterday. Nevertheless some carrots, basils, radishes and parsleys could be harvested soon. Our crew brought bean and tomato seeds to the GreenHab for one of our experiment. We soaked them into water during 24h and we will sow those seeds tomorrow.

Harvest: None

Support/supplies needed: None

Journalist Report – April 23rd

Today was our second simulation day. We woke up at 7:30 am and we had breakfast at 8:00. Yesterday, Carl made jam, we ate it, it was delicious. He tried to do choco pasta but for me it was a failure but the rest of the crew really appreciated it. A part of the crew had to go in EVA, but they had a problem with the battery of one of the rovers. The EVA was reported to afternoon.

At 9 am, each of us worked on his experience. Eleonore and Maxime did pictures of the sun with the solar telescope.

At 12 am, Chloé and Carl prepared the lunch. We ate quinoa with beans and corn. After meal, we discussed about Nathan’s experience and his martian constitution. We spoke about the political regime to introduce on Mars.

At 2pm, Eléonore and Maxime baked bread and today it was a success. Julien and Chloé cooked an apple cake, it smells tasty. After that, Maxime, Benjamin, Eléonore and Nathan went to EVA. The rover Opportunity had a battery problem. We had to give up it on the road.

At 6 pm, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. After the Cap Com, we will eat a “Gratin dauphinois” with beschamel sauce and pork sausages or lentils for vegetarians. And we will eat the apple cake for dessert.

Like usual, we will do a crew meeting to speak about the day.

EVA Report – April 23rd

Crew 211 EVA Report 23-04-2019

EVA #2

Author: Maxime Bernard

Purpose of EVA: Explore the surroundings in order to find a shelter in case of strong solar winds during another EVA.
Start time:16:00 pm
End time:18:00 pm

Narrative: The aim of this EVA was to explore Cando Chasma in order to find a shelter in case of strong solar winds. Unfortunately no cavity that could be used as shelter could be found. Moreover, we had to let one of the Rover on the road because of a battery failure (position: intersection of Stream Bed connector and Cactus road).

Destination: Candor Chasma

Coordinates (use UTM NAD27 CONUS): Latitude: 4251000, longitude: 521000

Participants: Eléonore Lieffrig (commander of the EVA), Benjamin Flasse (medical officer), Maxime Bernard, Nathan Pechon.
Road(s) and routes per MDRS Map: Cow Dung Road 0110 (2 kilometers towards the west), then Galileo road to reach cactus road. We finally drove half a kilometer towards the east. We then parked the rovers at panel 1104.

Mode of travel: Driving and hiking

Astronomy Report – April 23rd

Astronomy Report

Name: Lieffrig Eléonore

Crew: 211
Date: 23 April 2019


Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: None

Images submitted with this report: None

Problems Encountered: None


Solar Features Observed: Granules, prominences at the surface of the sun.

Problems Encountered: None

Sol Summary – April 23rd

Crew 211 Sol 2 Summary Report 23-APR-2019

Sol: 2

Summary Title: Feeling like a true Martian

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After the daily check-up performed by the HSO as part of his experiment, we enjoy a nice breakfast together. An EVA was planned at 9 am but as the Curiosity rover was not fully charged, we decided to postpone it. Most of the crew members worked therefore on their respective experiments while the crew astronomer and crew scientist got their first shots of the sun. We then worked altogether during the noon break on the writing of a Martian constitution with the crew journalist. During the early afternoon, some crew members prepared the first successful Martian bread. Half of the crew left at 4 pm for the planned EVA to Candor Chasma to look for possible shelters in case of high solar activity.

Look Ahead Plan: An EVA is planned in the morning to get back Opportunity near Candor Chasma and take additional soil samples. Half of the crew will stay at the base to work on their experiments. The crew astronomer will continue to take pictures of the sun in the morning and schedule a few observations with the Celestron telescope to get images from quasars. This will allow her to infer their luminosity and determine the evolution of their activity. The rest of the day will be devoted to the experiments.

Anomalies in work: No issues.

Weather: Hot, bright and sunny all day.

Crew Physical Status: Good, yesterday’s twisted ankle is less painful.

EVA: The aim of this second EVA was to explore the Candor Chasma region and look for caves. Underground structures could indeed be very helpful to protect the crew from solar wind/radiation in the case of increasing solar activity and CMEs (Coronal Mass Ejections). Unfortunately no caves have been found and we lost Opportunity near Candor Chasma. We’ll try to get the rover back during tomorrow’s EVA.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Astronomer Report, Operations Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site